A marketing analyst runs a bivariate regression model to forecast sales at a retail store based on advertising. Her model returns the following parameter estimates:

b0: 102

b1: 97.1

Standard Error: 70.95

She wished to forecast sales for a store where advertising equals 184. What is the forecasted standard deviation of the sales distribution for this store? Enter your answer rounded to 2 decimal places
Answer #1

The model is given by y = 102 + 97.1 x

if (x) advestising =184

y (sales) =102 + 97.1 * 184 =17968.4

Standard deviation of sales = b1*sx =97.1*70.95 = 6889.24